Mrs.Santos invests P21,000 at 8% interest rate.How much more must he invest at 9 1/2% interest to have an annual income of 9% of his total investments? (answer is P42,000 show your solution)
Answer by mananth(16946)

Mrs.Santos invests P21,000 at 8% interest rate.
let her invest x at 9 1/2% interest
annual income of 9% of his total investments

total investment =(x+21000)
8%*21000+9.5%x= 9%(x+21000)
multiply by 100
8*21000 +9.5x = 9x+189000
9.5x-9x= 189000-168000
0.5x=21000
multiply by 2
x=42000


(answer is P42,000 show your solution)